# Restoring Accessibility During Urban Rail Disruptions via Bus Network Redesign

### v1

@@ 3.5 Considerations about Computational Complexity
Our problem can be viewed as a complex combination of the Vehicle Routing Problem (VRP) [28] and the Capacitated Facility Location Problem (CFLP) [29], with an even more sophisticated objective function. Constraints (15) - (19) resemble those in VRP, handling vehicle allocation and line selection, while constraints (4) - (7) and parts of constraints (8) - (14) mirror CFLP constraints, addressing consolidation node selection, disrupted rail station assignment, and capacity limitations. Our objective function (Eq. 3) is more complex than those in VRP or CFLP; additionally, constraints (8) - (14) incorporate line design elements, further increasing the problem’s complexity. Given that both VRP and CFLP are known NP-hard problems, and our problem not only combines these two sub-problems but also adds extra constraints and a more intricate objective function, we can reasonably conclude that our problem is NP-hard.
@@ 4 A Two-stage Heuristic Algorithm
4:     for each regular line $l\in\pazocal{L}_{\text{REG}}$ do
5:        (Refer to Fig. 2) Compute extended line $l_{k}^{EX,a}$, originating from terminal $t^{a}_{l}$ of line $l$ and traversing all consolidation nodes $n$ within cluster $k$. To compute line $l_{k}^{EX,a}$, just apply the shortest path among nodes $n$ in cluster $k$. Perform the same computation to find extended line line $l_{k}^{EX,b}$.
@@
7:
8:  end for
9:  Stage 2: Bus Routes and Vehicle Allocation Optimization
10:  Decide which line should be connected to which cluster via line extensions the number of buses in each line (regular and extended) via the procedure explained in subsection “Stage 2”.
11:  return  Optimized redesigned PT network
To address the NP-hard problem of recovering accessibility during urban rail transit (URT) disruptions, we propose a two-stage heuristic algorithm in Alg. 1: in stage 1, we employ spatial clustering to group consolidation nodes into a reduced set of clusters, and we calculate optimal routing within each cluster; in stage 2, each line is associated to a cluster by one of its bus terminals, then extends through this terminal and all consolidation nodes of the associated cluster, following the routing decided in stage 1.
@@
Table 2: Table of cluster-based notation.
@@
$d_{l_{k}^{\text{EX},a}}~{};d_{l_{k}^{\text{EX},b}}$ The shortest distance of the hypothetical extended lines traversing the consolidation nodes within cluster $k$, $\forall l\in\pazocal{L}_{\text{REG}},\forall k\in\pazocal{K}$.
@@
$acc_{l}^{k}$ The accessibility of extended line $l$ to serve cluster $k$, $\forall l\in\pazocal{L}_{\text{EX}},\forall k\in\pazocal{K}$.
@@
$y_{lk}\in{0,1}$ 1 if extended line $l$ serves cluster k, 0 otherwise, $\forall l\in\pazocal{L}_{\text{EX}},\forall k\in\pazocal{K}$.
@@ 4.1 Stage 1: Demand-Driven Consolidation and Route Optimization
The first stage involves (i) associating disrupted rail stations to consolidation nodes based on constraint (4) - (7), (ii) clustering consolidation nodes, (iii) optimizing the traversal order for the extended lines of the suggested consolidation nodes of cluster $k$ to serve as an extended line $l$. To simplify, we associate rail station $d$ with consolidation node $n_{d}$, chosen as follows: if there exist consolidation nodes in $\pazocal{S}$ within distance $D_{max}$ from $d$, we set $n_{d}$ as the closest among them; otherwise, $n_{d}=d$, i.e. the consolidation node associated to the disrupted rail station is the station itself. We preferentially choose existing bus stops in $\pazocal{S}$ to consolidate disrupted demand in order to exploit the facilities available there (e.g., shelters, dynamic indication screens). Moreover, passengers can also board regular lines once they arrive at an existing bus stop. Therefore,
@@
In order for neighboring consolidation nodes to be served by the same bus routes and minimize the number of extended bus routes to avoid more impact on the regular line, we employ DBSCAN for clustering due to its ability to autonomously determine the number of clusters [30]. For each resulting cluster $k\in\pazocal{K}$ and for any regular line $l\in\pazocal L$, we compute a candidate extension line $l_{k}^{\text{EX},a}$, which represents the exact path that would be followed by extended line $l^{\text{EX},a}$ if it were associated to cluster $k$. To calculate this path, we formulate an open path problem based on constraints (8) - (14), calculating the minimum distance to traverse the consolidation nodes $n\in\pazocal{N}$ within cluster $k$. We do the same for extended line $l^{\text{EX},b}$. We repeat this calculation for each extension of all regular lines. Which extended line will be actually activated and traversing which cluster will be decided in Stage 2. We denote the length of hypothetical lines $l_{k}^{\text{EX},a},l_{k}^{\text{EX},b}$ by $d_{l_{k}^{\text{EX},a}},d_{l_{k}^{\text{EX},b}}$, respectively.
@@ 4.2 Stage 2: Bus Route Extension and Vehicle Allocation Optimization
In stage 2, we decide which extended line to associate to which cluster via line extension (variable $y_{lk},\forall k\in\pazocal{K}$) and show how many buses per unit of time we should allocate to each extended line and regular line (variable $x_{l}$, $\forall l\in\pazocal{L}=\pazocal{L}_{\text{REG}}\cup\pazocal{L}_{\text{EX}}$). Such decisions should be taken in order to maximize the objective (Eq. 3).
@@
For this reason, we replace the objective in Eq. 3 with a surrogate objective $f=f_{1}-f_{2}$, where $f_{1}$ is a proxy of overall accessibility (Eq. 3) and $f_{2}$ is the distance covered by extension buses. To calculate $f_{1}$, instead of computing the accessibility of all centroids and for each alternative, PT re-designed the configuration; we pre-computed the accessibility from stops and consolidation nodes in order to avoid combinatorial explosion. We associate regular bus line $l$, the following accessibility score:
$\text{acc}_{l}=\frac{1}{|\pazocal{S}_{l}|}\sum_{s\in\pazocal{S}_{l}}\sum_{c_{j}\in\pazocal{C},c_{i}\neq s}\frac{O_{c_{j}}}{T_{\pazocal{G}^{\text{DISR}}}(s,c_{j})},$ $\forall l\in\pazocal L_{\text{REG}}$ (21)
where $S_{l}$ is the set of all stations on regular line $l$, $s$ is a stop on the regular line $l$, $O_{c_{j}}$ is the number of opportunities within the tile with centroid $c_{j}$, $T{G}^{\text{DISR}}(s,c_{j})$ is the travel time from stops $s$ to centroid $c_{j}$ in the PT network structure during disruption; with slight abuse of notation, $c_{j}\neq s$ means that $c_{j}$ and $s$ must not be in the same tile. Similarly, we can get the accessibility of consolidation node $n$ as $acc_{n}$, as shown in Eq. 22; where again $c_{j}\neq n$, indicates that $c_{j}$ and $n$ must not be in the same tile. We associate to cluster $k$ the accessibility score as shown in Eq. 22.
@@
We finally associate an accessibility score with the hypothesis that an extension of the line $l\in l_{\text{REG}}$ will serve cluster $k\in\pazocal{K}$:
@@ 4.2.1 $f_{1}$ is defined as:
$f_{1}=\sum_{l\in\pazocal{L}_{\text{REG}}}acc_{l}\cdot x_{l}+\sum_{l\in\pazocal{L}_{\text{REG}}}\sum_{k\in\pazocal{K}}acc_{l}^{k}\cdot(x_{l^{EX,a}}+x_{l^{EX,b}})$ (25)
@@
  • More buses are allocated to lines with stops with high accessibility
  • More buses are allocated to extended lines that serve consolidation nodes with high accessibility
The rationale behind this choice is that if passengers can easily reach points with high opportunities, then, from there, they can reach more opportunities.
@@ 4.2.2 $f_{2}$ is defined as:
$f_{2}=\sum_{l\in\pazocal{L}_{\text{REG}}}\sum_{k\in\pazocal{K}}(d_{l^{EX,a}_{k}}\cdot x_{l^{EX,a}}+d_{l^{EX,b}_{k}}\cdot x_{l^{EX,b}})$ (26)
@@
[…]
$q_{k}\leq cap\cdot\sum_{l\in L_{EX}}x_{l}\cdot y_{lk},$ $\forall k\in\pazocal{K}$ (28)
$\sum_{l\in\pazocal{L}_{EX}}y_{lk}=1,$ $\forall k\in\pazocal{K}$ (29)
$x_{l^{EX,a}}\geq y_{l^{EX,a}k}~{};x_{l^{EX,b}}\geq y_{l^{EX,b}k}$ $\forall l\in\pazocal{L}_{\text{REG}}$ (30)
$x_{l^{EX,a}}\leq M\cdot y_{l^{EX,a}k}~{};x_{l^{EX,b}}\leq M\cdot y_{l^{EX,b}k}$ $\forall l\in\pazocal{L}_{\text{REG}}$ (31)
After applying clustering, adapted from constraint (15) - (19), constraint (28) ensures that the vehicle capacity for each cluster is not exceeded, constraints (29) ensures each extended line is assigned to serve exactly one cluster; while constraint. (30) and (31) ensure vehicles are only assigned in the event that an extended line is created;
@@ 5 Numerical Results
The performance of the methodology proposed herein is tested with numerical experiments using 2 case studies in two real cities, namely Évry-Courcouronnes, France, and Choisy-le-Roi, France. The integer programming model is coded and solved in Python using the IBM ILOG CPLEX 12.8 solver, running on an Intel Core i5 PC with 4.6 GHz speed and 16.0 GB of RAM. All solutions computed in this paper took several seconds to compute. The code is available at: this GitHub repository.
@@
Fleet size Cycle $\frac{\text{Cycle time}}{\text{Headway}}$ Eq. 4.4 of [25]
Average waiting time $\frac{headway}{2}$ [26]
@@
$D_{max}$ ((7)) 500 m -
Passenger capacity per bus (cap in (17)) 120 Passengers / Bus [31]
@@
          Optimal neighbourhood size $eps$ = 2 km -
@@
(a) Map of Évry-Courcouronnes created with Open-Street Map and folium [36]
@@
Figure 3: Évry-Courcouronnes’s public transport network and accessibility heat-map.
@@ 5.1 Disruption Case Study
We consider two circles of a 15 km radius centered in Évry-Courcouronnes and Choisy-le-Roi. We consider PT lines operating during the day (excluding night). As illustrated in Fig. 3(a), Évry-Courcouronnes PT consists of one urban rail line (RER D), one Tramway (T12), and 19 bus lines. Fig. 3(b) depicts the heat map of accessibility of accessibility (Eq. 3) where darker colors indicate higher accessibility. A similar analysis is performed for Choisy-le-Roi. This area is served by one Tramway (T9), one urban rail line (RER C), one metro line (M8), and 16 bus lines.
In our numerical experiment, we simulated complete disruptions of the rail lines (RER D in Évry-Courcouronnes and RER C in Choisy-le-Roi) lasting several hours, resulting in the closure of all rail stations within the affected areas, dramatically reducing accessibility. To ensure accurate distance measurements for each bus line, we utilized Open-Street Map data extracted via OSMnx [37]. To compute accessibility ((1)), we need first to count the number of opportunities inside each tile. The opportunities we consider are the amenities extracted from the open street map [38]. For simplicity, we extract all amenities without discerning among categories.

### v2

@@ 3.5 Considerations about Computational Complexity
Our problem can be viewed as a complex combination of the Vehicle Routing Problem (VRP) [29] and the Capacitated Facility Location Problem (CFLP) [30], with an even more sophisticated objective function. Constraints Eq. 15 - Eq. 19 resemble those in VRP, handling vehicle allocation and line selection, while constraints Eq. 4 - Eq. 7 and parts of constraints Eq. 8 - Eq. 14 mirror CFLP constraints, addressing consolidation node selection, disrupted rail station assignment, and capacity limitations. Our objective function (Eq. 3) is more complex than those in VRP or CFLP; additionally, constraints Eq. 8 - Eq. 14 incorporate line design elements, further increasing the problem’s complexity. Given that both VRP and CFLP are known NP-hard problems, and our problem not only combines these two sub-problems but also adds extra constraints and a more intricate objective function, we can reasonably conclude that our problem is NP-hard.
@@ 4 A Two-stage Heuristic Algorithm
4:  for each regular line $\ell\in\pazocal{L}_{\text{REG}}$ do
5:   (Refer to Fig. 2) Compute extended line $\ell_{k}^{EX,a}$, originating from terminal $t^{a}_{\ell}$ of line $\ell$ and traversing all consolidation nodes $n$ within cluster $k$. To compute line $\ell_{k}^{EX,a}$, just apply the shortest path among nodes $n$ in cluster $k$. Perform the same computation to find extended line line $\ell_{k}^{EX,b}$.
@@
7:end for
8:Stage 2: Bus Routes and Vehicle Allocation Optimization
9: Decide which line should be connected to which cluster via line extensions the number of buses in each line (regular and extended) via the procedure explained in subsection “Stage 2”.
10:return Optimized redesigned PT network
To address the computational complexity of our NP-hard optimization problem, we propose a two-stage heuristic algorithm that decomposes the computational burden while preserving problem structure integrity. In stage 1, we employ spatial clustering to group consolidation nodes into a reduced set of clusters and calculate optimal routing within each cluster. In stage 2, each line is associated to a cluster by one of its bus terminals, then extends through this terminal and all consolidation nodes of the associated cluster, following the routing decided in stage 1.
@@
Table 2: Notation for Cluster-Based Formulation
@@
$d_{\ell_{k}^{\text{EX},a}}~;d_{\ell_{k}^{\text{EX},b}}$ The shortest distance of the hypothetical extended lines traversing the consolidation nodes within cluster $k$, $\forall\ell\in\pazocal{L}_{\text{REG}},\forall k\in\pazocal{K}$.
@@
$acc_{\ell}$ Proxy measure of line $\ell$’s importance for accessibility, $\forall\ell\in\pazocal{L}_{\text{REG}}$
$acc_{n}$ Proxy measure of node $n$’s importance for accessibility, $\forall n\in\pazocal{N}$
$acc^{k}$ Proxy measure of cluster $k$’s importance for accessibility, $\forall k\in\pazocal{K}$
$acc_{\ell}^{k}$ Proxy measure of extended line $\ell$’s importance for accessibility when serving cluster $k$, $\forall\ell\in\pazocal{L}_{\text{REG}},\forall k\in\pazocal{K}$
@@
$y_{\ell k}\in{0,1}$ 1 if extended line $\ell$ serves cluster k, 0 otherwise, $\forall\ell\in\pazocal{L}_{\text{EX}},\forall k\in\pazocal{K}$.
@@ 4.1 Stage 1: Demand-Driven Consolidation and Route Optimization
The first stage involves (i) associating disrupted rail stations to consolidation nodes based on constraint Eq. 4 - Eq. 7, (ii) clustering consolidation nodes, (iii) optimizing the traversal order for the extended lines of the suggested consolidation nodes of cluster $k$ to serve as an extended line $\ell$. To simplify, we associate rail station $d$ with consolidation node $n_{d}$, chosen as follows: if there exist consolidation nodes in $\pazocal{S}$ within distance $D_{max}$ from $d$, we set $n_{d}$ as the closest among them; otherwise, $n_{d}=d$, i.e. the consolidation node associated to the disrupted rail station is the station itself. We preferentially choose existing bus stops in $\pazocal{S}$ to consolidate disrupted demand in order to exploit the facilities available there (e.g., shelters, dynamic indication screens). Moreover, passengers can also board regular lines once they arrive at an existing bus stop. Therefore,
@@
In order for neighboring consolidation nodes to be served by the same bus routes and minimize the number of extended bus routes to avoid more impact on the regular line, we employ DBSCAN for clustering due to its ability to autonomously determine the number of clusters [31]. For each resulting cluster $k\in\pazocal{K}$ and for any regular line $\ell\in\pazocal L$, we compute a candidate extension line $\ell_{k}^{\text{EX},a}$, which represents the exact path that would be followed by extended line $\ell^{\text{EX},a}$ if it were associated to cluster $k$. To calculate this path, we formulate an open path problem based on constraints Eq. 8 - Eq. 14, calculating the minimum distance to traverse the consolidation nodes $n\in\pazocal{N}$ within cluster $k$. We do the same for extended line $\ell^{\text{EX},b}$. We repeat this calculation for each extension of all regular lines. Which extended line will be actually activated and traversing which cluster will be decided in Stage 2. We denote the length of hypothetical lines $\ell_{k}^{\text{EX},a},\ell_{k}^{\text{EX},b}$ by $d_{\ell_{k}^{\text{EX},a}},d_{\ell_{k}^{\text{EX},b}}$, respectively.
@@ 4.2 Stage 2: Bus Route Extension and Vehicle Allocation Optimization
In stage 2, we decide which extended line to associate to which cluster via line extension (variable $y_{\ell k},\forall k\in\pazocal{K}$) and show how many buses per unit of time we should allocate to each extended line and regular line (variable $x_{\ell}$, $\forall\ell\in\pazocal{L}=\pazocal{L}_{\text{REG}}\cup\pazocal{L}_{\text{EX}}$). Such decisions should be taken in order to maximize the objective (Eq. 3).
@@
For this reason, we replace the objective in Eq. 3 with a surrogate objective $f=f_{1}-f_{2}$, where $f_{1}$ is a proxy of overall accessibility (Eq. 3) and $f_{2}$ is the distance covered by extension buses. Our two-stage heuristic algorithm effectively reduces computational complexity where only $acc_{\ell}$ (proxy measure of regular bus line $\ell$’s importance for accessibility, Eq. 21) and $acc_{n}$ (proxy measure of consolidation node $n$’s importance for accessibility, Eq. 22)) need to be pre-computed as input data. These accessibility scores are used directly in the calculation of our surrogate objective function Eq. 25. Once these metrics are available, our solution can be computed within seconds, making it suitable for responding to disruptions occurring on a timescale of hours. To calculate $f_{1}$, instead of computing the accessibility of all centroids and for each alternative, PT re-designed the configuration; we pre-computed the accessibility from stops and consolidation nodes in order to avoid combinatorial explosion. We associate regular bus line $\ell$ with the following accessibility proxy score:
$\text{acc}_{\ell}=\frac{1}{|\pazocal{S}_{\ell}|}\sum_{s\in\pazocal{S}_{\ell}}\sum_{c_{j}\in\pazocal{C},c_{i}\neq s}\frac{O_{c_{j}}}{T_{\pazocal{G}^{\text{DISR}}}(s,c_{j})},$ $\forall l\in\pazocal L_{\text{REG}}$ (21)
where $S_{\ell}$ is the set of all stations on regular line $\ell$, $s$ is a stop on the regular line $\ell$, $O_{c_{j}}$ is the number of opportunities within the tile with centroid $c_{j}$, note that $O_{c_{j}}$ is an input parameter, exogenously determined (i.e., we do not have control on it)”. $T{G}^{\text{DISR}}(s,c_{j})$ is the travel time from stops $s$ to centroid $c_{j}$ in the PT network structure during disruption; with slight abuse of notation, $c_{j}\neq s$ means that $c_{j}$ and $s$ must not be in the same tile. Similarly, we can get the accessibility proxy of consolidation node $n$ as $acc_{n}$, as shown in Eq. 22; where again $c_{j}\neq n$, indicates that $c_{j}$ and $n$ must not be in the same tile. We associate to cluster $k$ the accessibility proxy score as shown in Eq. 23.
@@
We finally associate an accessibility score with the hypothesis that an extension of the line $\ell\in l_{\text{REG}}$ will serve cluster $k\in\pazocal{K}$:
@@ 4.2.1 $f_{1}$ is defined as:
$f_{1}=\sum_{\ell\in\pazocal{L}_{\text{REG}}}acc_{\ell}\cdot x_{\ell}+\sum_{\ell\in\pazocal{L}_{\text{REG}}}\sum_{k\in\pazocal{K}}acc_{\ell}^{k}\cdot(x_{\ell^{EX,a}}+x_{\ell^{EX,b}})$ (25)
@@
  • More buses are allocated to lines with stops that are more important for accessibility
  • More buses are allocated to extended lines that serve consolidation nodes that are more important for accessibility
The rationale behind this choice is that by prioritizing lines and nodes with higher accessibility importance scores, we can more effectively restore the overall accessibility of the disrupted network.
@@ 4.2.2 $f_{2}$ is defined as:
$f_{2}=\sum_{\ell\in\pazocal{L}_{\text{REG}}}\sum_{k\in\pazocal{K}}(d_{\ell^{EX,a}_{k}}\cdot x_{\ell^{EX,a}}+d_{\ell^{EX,b}_{k}}\cdot x_{\ell^{EX,b}})$ (26)
@@
Constraint Eq.18
Constraint Eq.19
$q_{k}\leq cap\cdot\sum_{l\in L_{EX}}x_{\ell}\cdot y_{\ell k},$ $\forall k\in\pazocal{K}$ (28)
$\sum_{\ell\in\pazocal{L}_{\text{EX}}}y_{\ell k}=1,$ $\forall k\in\pazocal{K}$ (29)
$x_{\ell^{EX,a}}\geq y_{\ell^{EX,a}k}~;x_{\ell^{EX,b}}\geq y_{\ell^{EX,b}k}$ $\forall\ell\in\pazocal{L}_{\text{REG}}$ (30)
$x_{\ell^{EX,a}}\leq M\cdot y_{\ell^{EX,a}k}~;x_{\ell^{EX,b}}\leq M\cdot y_{\ell^{EX,b}k}$ $\forall\ell\in\pazocal{L}_{\text{REG}}$ (31)
After applying clustering, adapted from constraint Eq. 15 - Eq. 19, constraint Eq. 28 ensures that the vehicle capacity for each cluster is not exceeded, constraints Eq. 29 ensures each extended line is assigned to serve exactly one cluster; while constraint. Eq. 30 and Eq. 31 ensure vehicles are only assigned in the event that an extended line is created.
@@ 5 Numerical Results
The performance of the methodology proposed herein is tested with numerical experiments using 2 case studies in two real cities, namely Évry-Courcouronnes, France, and Choisy-le-Roi, France. The integer programming model is coded and solved in Python using the IBM ILOG CPLEX 12.8 solver, running on an Intel Core i5 PC with 4.6 GHz speed and 16.0 GB of RAM. All solutions computed in this paper took several seconds to compute.
@@
  Fleet size   $\dfrac{\text{Cycle time}}{\text{Headway}}$    Eq. 4.4 of [26]
  Average waiting time   $\dfrac{headway}{2}$    [27]
@@
  $D_{max}$ (Eq. 7)   500 m   -
  Passenger capacity per bus (cap in Eq. 17)   120 Passengers / Bus    [32]
@@
          Optimal neighbourhood size   epseps = 2 km   -
@@
(a) Map of Évry-Courcouronnes created with Open-Street Map
@@
Figure 3: Évry-Courcouronnes’s public transport network and accessibility heat-map. Symbols as in Fig. 1: circles = rail/tram stations, squares = bus stops, triangles = bus terminals.
@@ 5.1 Disruption Case Study
We consider two study areas defined as circles with a 15 km radius centered in Évry-Courcouronnes and Choisy-le-Roi, respectively. Within these areas, we analyze public transport lines operating during daytime hours, excluding night services. As illustrated in Fig. 3(a), Évry-Courcouronnes PT consists of one urban rail line (RER D), one Tramway (T12), and 19 bus lines. In this figure, the different types of transport lines and stations are represented using the same symbol convention as in Fig. 1. Rail and tram stations (collectively referred to as rail transit stations) are depicted as circles, bus stops as squares, and bus terminals as triangles throughout all figures in the paper. Fig. 3(b) depicts the heat map of accessibility of accessibility (Eq. 3) where darker colors indicate higher accessibility. A similar analysis is performed for Choisy-le-Roi. This area is served by one Tramway (T9), one urban rail line (RER C), one metro line (M8), and 16 bus lines.
In our numerical experiment, we simulated complete disruptions of the rail lines (RER D in Évry-Courcouronnes and RER C in Choisy-le-Roi) lasting several hours, resulting in the closure of all rail stations within the affected areas, dramatically reducing accessibility. To ensure accurate distance measurements for each bus line, we utilized Open-Street Map data extracted via OSMnx [37]. To compute accessibility (xwEq. 1), we need first to count the number of opportunities inside each tile. The opportunities we consider are the amenities extracted from the open street map [38]. For simplicity, we extract all amenities without discerning among categories.